What is the concentration of HCL, when 50.0 mL of the solution was neutralized by 12 mL of a 0.35 mol/L solution of KOH?
1
Expert's answer
2012-12-25T09:54:25-0500
For answering this question you need to use next law:
N1V1 = N2V2,
where N is normality (in this case it is the same that molaritiy) and V is volume,

So N2 = N1V1/V2 = 12*0.35 / 50.0 = 0.084 mol/l